OpenStudy (anonymous):

an objects motion is describe by the equation d=3sin(8pi t)-2. the displacement d is measured in meter. the time t is measured in seconds. a. what is the object position (in meter)at t=0? b. what is the object maximum displacement(in meter) from its t=0 position? c. how much time(in sec) is required for one oscillation d. what is the frequency(measured in Hz) of this oscillation

OpenStudy (anonymous):

in this equation \[\omega=8\pi\] a) d=0 b)3 c)\[\theta=\omega t\] for 1 oscilation t=T & theta=2pi so it yields T d)\[\omega=2\pi \nu\]